OpenStudy (anonymous):

This question came up in my math test and I don't know how to solve it! Please help! For the quadratic equation, x^2 - 4x + c = 0 where the equation has only one solution... Find the value of c.

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x=\frac{4\pm\sqrt{16-4c}}{2}\]\[x=2\pm\sqrt{4-c}\]

OpenStudy (anonymous):

Ok

OpenStudy (unklerhaukus):

there will be only one solution when \(x=2\pm0\)

OpenStudy (anonymous):

Alright I get it thanks

hero (hero):

A quadratic equation will have one solution when when and only when the discriminant equals zero. Therefore \(b^2 - 4ac = 0\) a = 1 b = -4 c = c \((-4)^2 - 4(1)c = 0 \\ 16 - 4c = 0 \\ 16 = 4c \\ c = 4\)

hero (hero):

@jadecat